OpenStudy (tylermckinney16):

Calculate the cost of the trip: Gas: 400 miles divided by 30 miles per gallon times $2.75 per gallon Lodging: Cheap and free campgrounds with a tent. Stretching trip over five days. $75 Food: Slightly cheaper from cooking meals at the campground. $350 TOTAL: ?

OpenStudy (chiunlee12):

okay, so all you need to do is add everything. So, if we simplify all this wording into equations, Gas= 400/30 x 2.75 Lodging= 75 Food= 350 And all you need to do is to add all these values together, so: Gas + Lodging + Food = TOTAL

OpenStudy (danjs):

\[gas ~~cost~~=\large \frac{ 400~mi }{ 1 }*\frac{ 1~~gal }{ 30~~mi }*\frac{ $2.75 }{ 1~~gal }=\]

OpenStudy (tylermckinney16):

$461.6

OpenStudy (tylermckinney16):

that the answer?

OpenStudy (caponeishere):

No, OS closing is.

OpenStudy (tylermckinney16):

@chiunlee12

OpenStudy (chiunlee12):

I think it is

OpenStudy (chiunlee12):

461.67 to be more accurate haha